OpenStudy (rissyroo118):

Need help in accounting will give medals too each question I post please we and thank you

OpenStudy (will.h):

tag me when you do.

OpenStudy (rissyroo118):

Using the interest formula compute the interest and maturity values. #1 $4,000 is the principal Interest term is 11.5% Rate is 60 days @Will.H

OpenStudy (will.h):

you can use exponential function f(x)= p(1+r)^x where p= initial number r is the percent rage x is the time f(x)= 4000(1+ 11.5/100)^60

OpenStudy (will.h):

just note that t normally is years. but we can convert 60 days into years. let me do it. 1m

OpenStudy (will.h):

The formula above is used when finding intrest rates over the years. However. In your question the time is in days. there's another formula \[P(1+rt)\] where p s the initial number and t is time in days and is the percent so p=4000 and r= 11.5% which means 0.115 (after dividing on 100) and t = 60 therefore 4000(1+ 0.115(60)) which would equal 31600$ hope that helps

OpenStudy (rissyroo118):

So the interest would be after all $31,600?

OpenStudy (will.h):

that's the principal after 60 days at a growth of 11.5%

OpenStudy (will.h):

make sense?

OpenStudy (rissyroo118):

Hm not sure , I understand but thanks. I'm gonna try to use the info you gave.